# Stability of discrete-symmetry flocks: sandwich state, traveling domains and motility-induced pinning

### v1

@@ Appendix F Droplet pinning at low temperature
Fig. 17 demonstrates the fate of an inserted droplet within the polar-ordered phase at large $\beta$ and small $D$. Over time, the droplet freezes [Fig. 17(a)] with its front becoming jammed due to the accumulation of opposite-polarity particles $(t=64$ and 160), caused by the motility-induced interface pinning. This phenomenon is further explained by the magnetization profiles of the ordered state ($\sigma=3$) and the droplet state ($\sigma=1$) shown in Fig. 17(b), where the sharp opposite peaks of the two magnetization profiles indicate a high-density interface created by opposite polarity particles. Fig. 17(c) shows the position of the droplet front ($x_{f}$) over time, exhibiting that after an initial movement in the biased direction, the droplet front remains stationary. A similar pinning of a droplet has been observed in the active $p$-state clock model for $p\leq 4$ Woo and Noh (2024). However, in the $p=8$ state, where particles have more orientational degrees of freedom, droplet pinning does not occur Woo and Noh (2024).

### v2

@@ Appendix G Droplet pinning at low temperature
Fig. 19 demonstrates the fate of an inserted droplet within the polar-ordered phase at large $\beta$ and small $D$. Over time, the droplet freezes [Fig. 19(a)] with its front becoming jammed due to the accumulation of opposite-polarity particles $(t=64$ and 160), caused by the motility-induced interface pinning. This phenomenon is further explained by the magnetization profiles of the ordered state ($\sigma=3$) and the droplet state ($\sigma=1$) shown in Fig. 19(b), where the sharp opposite peaks of the two magnetization profiles indicate a high-density interface created by opposite polarity particles. Fig. 19(c) shows the position of the droplet front ($x_{f}$) over time, exhibiting that after an initial movement in the biased direction, the droplet front remains stationary. A similar pinning of a droplet has been observed in the active $p$-state clock model for $p\leq 4$ jd2024MIP. However, in the $p=8$ state, where particles have more orientational degrees of freedom, droplet pinning does not occur jd2024MIP.